OpenStudy (abbycross167):

Can someone please help me with this question? I have no clue how to do it.... Graph the inequality 102c + 150m + 41 > 500

OpenStudy (anonymous):

Ok so have you've worked with linear equations in standard form right? Like for ex. 3x+2y=10?

OpenStudy (abbycross167):

yes sir/ma'am

OpenStudy (anonymous):

Oh lol no need for the sir Anyway in the case of the example above you could graph it by finding the x and y intercepts,... how could you do that in 3x+2y=10?

OpenStudy (abbycross167):

so x=3 and y=5?

OpenStudy (anonymous):

Be careful with your x If you substitute 0 for y, which is what an x intercept is because it's y is 0, you get 3x+2(0)=10 3x+0=10 3x=10 x=3.333333333 repeating

OpenStudy (abbycross167):

oh ok

OpenStudy (anonymous):

Ok so do the same thing with the inequality, except make sure it is a DASHED line when you draw and your SHADING is on the correct side because this is an inequality

OpenStudy (anonymous):

I suggest moving the 41 to the other side before you start substituting 0's for x and y

OpenStudy (abbycross167):

what do you mean?

OpenStudy (anonymous):

Well you know how you substituted 0 to get the x and y intercepts in the example right? Do the same thing to your problem, I'm just saying move 41 to the right side to get c and m by themselves

OpenStudy (abbycross167):

So 41+102c + 150m > 500 ?

OpenStudy (anonymous):

Well what I mean by move the 41 over is to subtract 41 from both sides You know, getting rid of it on the left and doing the same thing to the right side

OpenStudy (anonymous):

So you would get 102c+150m+41-41>500-41

OpenStudy (anonymous):

What would that be?

OpenStudy (abbycross167):

102c+150>500 OR 102c+150>459?

OpenStudy (anonymous):

It's the second one right? Because you have to count the 41 Now substitute 0 both for m and c and find the m(x) and c(y) intercepts

OpenStudy (anonymous):

Actually that's an interesting question Does it specify if m is x or m is y or vice versa

OpenStudy (abbycross167):

no sir I don't believe so, would you like me to show you the whole question?

OpenStudy (anonymous):

Umm yeah sure

OpenStudy (anonymous):

I'm sorry it's taking me so long to reply I'm doing my homework while doing this and I'm getting distracted

OpenStudy (abbycross167):

it's no problem! Thank you kindly for helping me!!

OpenStudy (abbycross167):

"You want to plan a nutritious breakfast. It should supply at least 500 calories or more. Be sure your choices would provide a reasonable breakfast. Suppose you want to have cereal, milk, and one glass of tomato juice. Let c be the number of cups of cereal and m the number of cups of milk. The inequality 102c + 150m + 41 > 500 models the situation." Graph the inequality. (Assume you can eat partial eggs and drink partial apple juices.)

OpenStudy (anonymous):

Well anyway I guess it doesn't matter which is the x and which is the y So anyway now you have \[102c+150m>459\] Right? So now substitute 0 for c, What do you get for m? Substitute 0 for m, what do you get for c?

OpenStudy (anonymous):

just ignore the greater than for now, imagine it as an equal sign

OpenStudy (abbycross167):

So if c=0 and m=0 it would look like this? 102+150>459?

OpenStudy (anonymous):

well don't set both to 0 Do 1 at a time So do 102(0)+150m>459, solve for m Then start over and do 102c+150(0)>459, solve for c

OpenStudy (abbycross167):

So 102(0)+150m>459 m=150?

OpenStudy (anonymous):

Not quite Here let me show you 102(0) would just equal 0 So now you have 0+150m>459 Which is just 150m>459 So now you DIVIDE by 150 to get m>3.06

OpenStudy (anonymous):

Do you understand why you divide? You do so because you have \[150m>459\] Dividing by 150 could get you \[\frac{ 1 }{ 150 }\times \frac{ 150m }{ 1 }= 459\div150\]

OpenStudy (anonymous):

that equals 150m/150 which simplifies to just m

OpenStudy (abbycross167):

oh ok, so then what?

OpenStudy (anonymous):

You try the same with the second equation